OpenStudy (anonymous):

Each statement describes a transformation of the graph of y = x^2. Which statement correctly describes the graph of y = (x + 9)^2 + 4 ? A. It is the graph of y = x^2 translated 4 units up and 9 units to the left. B. It is the graph of y = x^2 translated 9 units up and 4 units to the left. C. It is the graph of y = x^2 translated 9 units up and 4 units to the right. D. It is the graph of y = x^2 translated 4 units up and 9 units to the right.

OpenStudy (zale101):

y=(x-h)^2+k h is horizontal shift, and in the equation (x-h)^2, the graph shifts to the right h units k is the vertical shift, in the equation +k, the graph vertically shifts up k units y = (x + 9)^2 + 4 y = (x -(- 9))^2 + 4 Now, what can you conclude about this?

OpenStudy (anonymous):

Theyre practically the same thing except more of a negative in the second then the first

OpenStudy (zale101):

y = (x -(- 9))^2 + 4 is the same thing as y = (x + 9)^2 + 4, yes and it means the graph shifts to the left or right, up or down?

OpenStudy (zale101):